A cake is removed from an oven at 325F and cools to 150F after 25 minutes in a 68F room. How long will it take to cool to 115F?
Answer by fractalier(2101)

I believe you need Newton's Law of Cooling for this one...
Let's see if I can remember this one...something like
T - Tr = (To - Tr)*e^(-kt)
now we plug in
150 - 68 = (325 - 68)*e(-25k)
and find k...
e^(-25k) = .319
-25k = -1.142
k = .0457
Now put that back in to the original equation with the new data and solve for t...
T - Tr = (To - Tr)*e^(-.0457t)
115 - 68 = (325 - 68)*e^(-.0457t)
e^(-.0457t) = 47/257
e^(-.0457t) = .1829
-.0457t = ln .1829 = -1.699
t = 37.2 minutes